# MoESD: Mixture of Experts Stable Diffusion to Mitigate Gender Bias

### v1

@@ Abstract
Text-to-image models are known to propagate social biases. For example when prompted to generate images of people in certain professions, these models tend to systematically generate specific genders or ethnicity. In this paper, we show that this bias is already present in the text encoder of the model and introduce a Mixture-of-Experts approach by identifying text-encoded bias in the latent space and then creating a bias-identification gate. More specifically, we propose MoESD (Mixture of Experts Stable Diffusion) with BiAs (Bias Adapters) to mitigate gender bias. We also demonstrate that a special token is essential during the mitigation process. With experiments focusing on gender bias, we demonstrate that our approach successfully mitigates gender bias while maintaining image quality.
MoESD: Mixture of Experts Stable Diffusion to Mitigate Gender Bias
Guorun Wang Imperial College London g.wang23@imperial.ac.uk                        Lucia Specia Imperial College London l.specia@imperial.ac.uk
@@ 1 Introduction
In recent years, large language and vision models such as ChatGPT 4 OpenAI (2023), DALL·E Ramesh et al. (2021) and Stable Diffusion Rombach et al. (2022); Podell et al. (2023) have ushered the era of AI generated content. However, research has shown that these generative models often exhibit social biases during the content generation process, especially in text-to-image generation. For instance, models tends to generate more images of man when provided with prompts like “a successful CEO”, and more images of women when provided with prompts like “a paralegal” Friedrich et al. (2023); Luccioni et al. (2023).
@@
In pre-processing methods, eliminating bias in the training corpus is a difficult challenge that offers no guarantees Hamidieh et al. (2023). For prompt-engineering, although leveraging specific prompts to instruct the model (e.g. “a photo of a female plummer”) Friedrich et al. (2023) can work to avoid biases, this is not how the average user prompts these models, do not a solution in practice. For changing model weights, the resource-intensive nature of re-training models poses challenges, requiring vast amounts of data (89k text-image pairs in Esposito et al. (2023)) and full fine-tuning of the model.
Our work identifies existing biases in pre-trained models and effectively mitigates them by parameter-efficient fine-tuning, which only requires a small amount of data (1.5K) and parameters (5.6%). Our contributions can be summarized as follows:
  • We measure the gender skew in text to assess gender bias in embeddings.
@@
  • We apply special tokens to aid the BiAs in better understanding the biased data and demonstrate that it is essential for mitigation.
  • We successfully mitigate the gender bias in Stable Diffusion while maintaining image quality.
@@ 2 Related Work
Biases in multimodal settings have attracted increasing attention. Several studies have evaluated multimodal models and discovered that they inherit and propagate many biases Agarwal et al. (2021); Cho et al. (2023). Prompt learning and engineering have been widely utilized in vision-language models and generative models Berg et al. (2022); Friedrich et al. (2023). Specifically, Fair Diffusion Friedrich et al. (2023) addresses biases through prompt engineering: users insert prompts to generate fair images with the assistance of specific guidance on the SEGA model Brack et al. (2023). Subsequently, image generation is guided towards a fairer outcome through manual semantic editing in the latent space of biased concepts.
Approaches to offset bias representation are also increasingly popular: Seth et al. (2023) employs additive residual image representations to mitigate biased representations, while Esposito et al. (2023) focuses on fine-tuning the model to achieve fairness. Chuang et al. (2023) proposed a method to project out biased directions in text embeddings to create fair generative models. This approach leverages positive pairs of prompts to debias embeddings effectively. By generating embeddings of prompts such as “a photo of a [class name] with [spurious attribute]”, a calibrated projection matrix, as shown in Equations 1, is optimized. After projection, the embedding should only contain information about the “[class name]” with no spurious information (e.g., gender).
Equation 1 illustrates the regularization of the difference between the projected embeddings of the set of positive pairs $S$, where $(z_{i},z_{j})$ represents the embedding of prompt pair $(i,j)$ in $S$, which describes the same class but with different spurious attributes (e.g., gender). The loss function encourages the linear projection $P$ to be invariant to the difference between the spurious attributes (details in Appendix A.4).
@@ 3 Method
In this section we describe our method. First, we explore the projection matrix from Chuang et al. (2023) (Equation 1) to assess inheriting gender biases in Stable Diffusion. Second, we introduce the Mixture of Experts (MoE) to Stable Diffusion and add fine-tuned BiAs (experts), which combine bias identification gates and special tokens to aid in understanding the biased data. This results in our MoESD-BiAs approach.
@@ 3.1 Identifying Biases in Stable Diffusion
Stable Diffusion, the text-to-image model in our work, generates images from text by switching diffusion process from pixel space to latent space to generate images. Given an image $x\in\mathbb{R}^{H\times W\times 3}$ in RGB space, the encoder $\mathcal{E}$ encodes $x$ into a latent representation $z=\mathcal{E}(x)$, and the decoder $D$ reconstructs the image from the latent, giving $x^{\prime}=D(z)=D(\mathcal{E}(x))$, where $z\in\mathbb{R}^{h\times w\times c}$. A domain-specific encoder $\tau_{\theta}$ projects $y$ (language prompt) to an intermediate representation $\tau_{\theta}(y)\in\mathbb{R}^{M\times d_{\tau}}$, which is then mapped to the intermediate layers of the U-Net via a cross-attention layer implementing to condition the latent $z$. Based on image-conditioning pairs, the conditional LDM is learned via Equation 2.
@@
However when employing prompts like “a photo of a [occupation]”, the model has been shown to exhibit significant gender bias with stereotypes Friedrich et al. (2023); Chuang et al. (2023), which manifests throughout the entire process.
While research efforts have been dedicated to addressing biases in the iterative image denoising process within U-Net Ronneberger et al. (2015), the convolutional network component of Stable Diffusion Esposito et al. (2023), our investigation has revealed a distinct bias emerging after the text embedding stage, when utilizing the text encoder of different versions of Stable Diffusion. This means that bias is amplified throughout the process.
Appendix A.2 shows a visualization of the dimensionality-reduced embeddings encoded by the text encoder of two versions of Stable Diffusion [^1] https://huggingface.co/runwayml/stable-diffusion-v1-5  [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1-base using T-SNE, with prompts like “The photo of the face of a [occupation]”, where the [occupation] are from the top 8 male-biased and top 8 female-biased occupations from Fair Diffusion Friedrich et al. (2023) [^3] https://github.com/ml-research/Fair-Diffusion/blob/main/results_fairface_generated_1-5.txt. There is a clear boundary between two gender-biases occupation encoded embeddings in both versions of Stable Diffusion. In other words, embeddings show gender biases regardless of whether they are encoded by CLIP or OpenCLIP Radford et al. (2021); Schuhmann et al. (2022); Cherti et al. (2023), indicating that the text encoder already contains gender biases, which then conditions the U-Net. Thus, it is not enough to mitigate biases in the latent space of the U-Net, we need to address the text-encoded bias as well.
@@ 3.2 Measuring Text-Encoded Bias
Although Chuang et al. (2023) claims great success in measuring bias in text embeddings and projecting it to achieve fairness, we have identified an issue. When attempting to switch the original prompt into another direction to achieve balance, the projection may take it into another direction that can introduce other biases. In the Appendix A.6, we present failure cases of gender mitigation using Chuang et al. (2023)’s method, which resulted in only male faces being generated.
@@
We conduct zero-shot and unsupervised classification on embeddings, using prompts like “The photo of the face of a [occupation]” and a pretrained Stable Diffusion text encoder. Our goal is to measure gender bias within these prompts and their embeddings. Altering the weights of the pretrained text encoder is risky, as it could lead to the model forgetting information that is not related to gender, which has learned from extensive data, so the Stable Diffusion text encoder is frozen. Another challenge is the lack of clear labels, as we only have prompts, embeddings, and statistics for each occupation generated by Fair Diffusion. These statistics are based on 250 images from the original Stable Diffusion and cannot be quantitatively analyzed or treated as a simple binary classification task due to the small sample size. Moreover, the embeddings are subject to change based on the hyperparameters of the pretrained model, meaning the statistics can only serve as a benchmark for testing our approach rather than for model learning. Therefore, our analysis relies solely on prompts and model weights.
As shown in Equation 1, $z$ is the prompt embedding, and $P$ is the projection process. Prompts $z_{0}$ applying Calibration Matrix can be written as $P^{*}z_{0}=P_{0}\left(I+\frac{\lambda}{\left|S\right|}\sum_{(i,j)\in S}(z_{i}-z_{j})(z_{i}-z_{j})^{T}\right)^{-1}z_{0}$, derived from Equation 1.
@@
$\Delta S(z_{0},z_{t},P_{0},P^{*})$ $=\text{Similarity}(P_{0}z_{0},P_{0}z_{t})$
$\quad-\text{Similarity}(P^{*}z_{0},P_{0}z_{t})$ (3)
@@
$\mathscr{G}(z_{0})$ $=\Delta S(z_{0},z_{male},P_{0},P^{*})$
$\quad-\Delta S(z_{0},z_{female},P_{0},P^{*})$ (4)
We assume the gender skew is male when $\mathscr{G}(z_{0})>0$ while the gender skew is female $\mathscr{G}(z_{0})<0$. If $\Delta S(z_{0},z_{t},P_{0},P^{*})$ is larger, then the disparity between two prompts-before and after applying the Calibration Matrix is also larger. This results in a greater divergence between the original embedding direction and the projected embeddings. Consequently, the strength of the redirected embeddings is increased, indicating a stronger gender preference before redirecting. In essence, if the redirection strength is substantial, it indicates a more forceful adjustment away from the original gender direction, which means higher gender skew.
@@ 3.3 Defining the Bias Identification Gate
We utilize the Pearson Correlation Coefficient to measure similarity and check statistical approximation of genders in Fair Diffusion occupations, with further details provided in Appendix A.7. If the male count for a particular occupation surpasses half of the total, we assume this occupation is male-skewed; otherwise, it is female-skewed. To assess the effectiveness of our approach when designing the Bias Identification Gate, we compare the frequency-based labels with our calculations derived from Equation 3.2 to compute accuracy. We achieve an accuracy rate of 79%, indicating that our bias measurement approximation, achieved through task reframing, aligns with both intuition and statistical trends.
The accuracy is lower when it comes to occupations such as insurance agents (details in Appendix A.9.1), as these are mostly likely at the boundary between male and female. In Table 1, we show mitigation of biases even when classifications of these challenging occupations are incorrect.
@@ 3.4 BiAs Experts: Bias Adapter Experts
The intuition is to guide the model to generate more female when the original embedding exhibits the male skew and vice versa. To achieve this, we divide our training text-image pairs into two groups: male and female. We then fine-tune the model separately to generate male-biased and female-biased experts. When the gate is activated, it guides the experts according to the following rules: male skew for female-biased experts and female skew for male-biased experts. This process is called bias fine-tuning.
Figure 1: The Architecture of BiAs. We apply BiAs to the cross-attention of the U-Net. In the figure, the left BiAs represents the female expert, the right one represents the male expert, and the middle represents the original cross-attention. The conditional information is processed by the Bias Identification Gate to determine which experts to choose - either male, female, or none of them. In the end, the chosen BiAs will process the input and be added together with the original cross-attention. The detailed BiAs architecture includes BiAs of the Q, K, V, and output matrices (for simplicity, we omit them in the figure).
Adapters Hu et al. (2022); Houlsby et al. (2019); Karimi Mahabadi et al. (2021) provide a method to freeze the model and introduce a new, trainable weight matrix, which significantly reduces both the time and memory required for training.
We incorporate this method into creating bias experts, making the bias fine-tuning parameter-efficient. We freeze the parameters in the U-Net and add adapters on the cross-attention layers of the U-Net, as illustrated in Figure 1. We initialize the adapters following the principles of LoRA Hu et al. (2022): the weight parameters of the first matrix $W_{\text{down}}^{\prime}$ are determined by a Gaussian function, while the parameters of the second matrix $W_{\text{up}}^{\prime}$ are initialized as a zero matrix. Consequently, the added pathway is initially zero during training and does not impact the result, resembling the original output.
Since the adapters are randomly initialized, they do not inherit biases from Stable Diffusion fine-tuning, which makes them more effective in achieving fairness, as detailed in Section 4.2.2. By using adapters as the bias experts, we achieve better fairness scores and reduce the trainable ratio to only 5.6%.
For bias fine-tuning, we generate biased images with specific male or female characteristics so that the bias expert will learn stereotypes from the data. We employ a technique from Dreambooth Ruiz et al. (2023), utilizing a special token to help BiAs to better understand the biased data. Dreambooth adds a special token, a unique identifier, to the prompt, and uses a few fine-tuning images of a subject as input to “personalize” the model. In our work, we use the special token to “personalize” the adapters of the biased information.
@@
For fine-tuning data, we generate our training image-text pairs using Stable Diffusion-XL with the following prompt: “A [gender] + [race] + [occupation].” The prompts generate 1530 images with different genders, races and occupations to ensure variety, which then fine-tune the model using the special token. The results demonstrate that the special token yields better performance than ordinary fine-tuning, as detailed in Section 4.2.2.
@@
What distinguishes our work from Fair Diffusion Friedrich et al. (2023) is that we do not manually engage in prompt engineering (human intervention), in other words, we do not modify the prompt. The special token we add can be any token. Moreover, even without the special token, our approach still mitigates gender bias.

### v2

@@ Abstract
Text-to-image models are known to propagate social biases. For example, when prompted to generate images of people in certain professions, these models tend to systematically generate specific genders or ethnicities. In this paper, we show that this bias is already present in the text encoder of the model and introduce a Mixture-of-Experts approach by identifying text-encoded bias in the latent space and then creating a Bias-Identification Gate mechanism. More specifically, we propose MoESD (Mixture of Experts Stable Diffusion) with BiAs (Bias Adapters) to mitigate gender bias in text-to-image models. We also demonstrate that introducing an arbitrary special token to the prompt is essential during the mitigation process. With experiments focusing on gender bias, we show that our approach successfully mitigates gender bias while maintaining image quality.
@@ 1 Introduction
In recent years, large language and vision models such as ChatGPT 4 (OpenAI 2023), DALL·E (Ramesh et al. 2021) and Stable Diffusion (Rombach et al. 2022; Podell et al. 2024) have ushered the era of AI-generated content. However, research has shown that these generative models often exhibit social biases during the content generation process, especially in text-to-image generation. For instance, models tend to generate more images of men when provided with prompts like “a successful CEO”, and more images of women when provided with prompts like “a paralegal” (Friedrich et al. 2023; Luccioni et al. 2023).
@@
In pre-processing methods, eliminating bias in the training corpus is a difficult challenge that offers no guarantees (Hamidieh et al. 2023).
For prompt-engineering, although leveraging specific gender attributes to instruct the model (e.g., “a photo of a female plumber”) can help avoid biases (Berg et al. 2022; Friedrich et al. 2023), it is not a practical solution as it does not fix the model and one cannot guarantee that the model will always be used with bias mitigating prompts.
For changing model weights, the resource-intensive nature of re-training models poses challenges, requiring vast amounts of data (89k text-image pairs in Esposito et al. (2023)) and fine-tuning of the model.
Although recent methods (Friedrich et al. 2023; Zhang et al. 2023; Parihar et al. 2024; Gandikota et al. 2024; Wang et al. 2024) demonstrate that explicitly specifying gender attributes in the prompt and token, or providing gender attribute guidance in the latent space during generation can yield good results, we aim to reduce this manual interventions to the lowest possible level.
We propose a method based on Mixture of Experts for Stable Diffusion with Bias Adapters, which identifies existing biases in pre-trained models and effectively mitigates them without requiring gender-guided prompts during generation. We modify the actual model weights, but in a parameter-efficient manner, requiring only a small amount of data (1.5K) and parameters (5.6%) to be updated. Our contributions can be summarized as follows:
  • We measure the gender skew in text to assess gender bias in embeddings and propose a Bias Identification Gate mechanism.
@@
  • We investigate the use of special tokens to aid the BiAs in better understanding the biased data and demonstrate that it is essential for mitigation, even for challenging professions.
  • We show that the proposed method successfully mitigates the gender bias in Stable Diffusion generation while maintaining image quality.
  • We conducted additional experiments on skin tone biases to show the generalisation ability of our method to other biases.
@@ 2 Related Work
Biases in multimodal settings have attracted increasing attention. Several studies have evaluated multimodal models and found that they inherit and propagate various biases (Agarwal et al. 2021; Cho, Zala, and Bansal 2023; Bansal et al. 2022). For example, Bansal et al. (2022) introduced the ENTIGEN framework to evaluate and mitigate gender bias. Prompt and token learning with attributes, as well as attribute editing in the latent space, have been widely utilized in vision-language models and generative models (Berg et al. 2022). Specifically, Fair Diffusion (Friedrich et al. 2023) explicitly inserts gender attributes into the SEGA model (Brack et al. 2023). ITI-GEN (Zhang et al. 2023), similar to Fair Diffusion, learns a set of prompt embeddings to represent gender attribute categories. Balancing Act (Parihar et al. 2024) trains a classifier to provide gender attribute guidance in the latent space. UCE (Gandikota et al. 2024) directly edits concepts based on phrases and attributes, while Concept Algebra (Wang et al. 2024) manipulates and combines concepts in the latent space using algebraic operations. Subsequently, image generation is guided towards a fairer outcome through manual semantic editing in the latent space of biased concepts. Although these methods can yield better fairness results in the generated images, our focus is on addressing the biases inherent in the model itself. Therefore, our method learns to modify the model itself and exploits arbitrary tokens and embeddings without gender guidance in the generation process.
Approaches to offset bias representation in the model are also increasingly popular: Seth, Hemani, and Agarwal (2023) employs additive residual image representations to mitigate biased representations, while Esposito et al. (2023) and Fairness Fine-tuning (Shen et al. 2024) focuses on fine-tuning the model to achieve fairness. DebiasVL (Chuang et al. 2023) proposed a method to project out biased directions in text embeddings to create fair generative models leveraging positive pairs of prompts to debias embeddings. By generating embeddings of prompts such as “a photo of a [class name] with [spurious attribute]”, a calibrated projection matrix, as shown in Equations 1, is optimized. After projection, the embedding should only contain information about the “[class name]” with no spurious information (e.g., gender).
Equation 1 illustrates the regularization of the difference between the projected embeddings of the set of positive pairs $S$, where $(z_{i},z_{j})$ represents the embedding of prompt pair $(i,j)$ in $S$, which describes the same class but with different spurious attributes (e.g., gender). The loss function encourages the linear projection $P$ to be invariant to the difference between the spurious attributes (details in Appendix D). $P_{0}$ is the encoding process of the pre-trained Stable Diffusion.
@@
In our work, we utilize the Projection $P$ from Equation 1 to calculate gender skew from the prompt and use the information to better fine-tune the model to offset bias.
@@ 3 Method
In this section, we describe our method. First, we explore the projection matrix from DebiasVL (Equation 1) to assess inheriting gender biases in Stable Diffusion. Second, we introduce the Mixture of Experts (MoE) to Stable Diffusion and add fine-tuned BiAs (experts), which combine bias identification gates and special tokens to aid in understanding the biased data. This results in our MoESD-BiAs approach.
@@ 3.1 Identifying Biases in Stable Diffusion
Stable Diffusion, the text-to-image model in our work, generates images from text by switching the diffusion process from pixel space to latent space to generate images. Given an image $x\in\mathbb{R}^{H\times W\times 3}$ in RGB space, the encoder $\mathcal{E}$ encodes $x$ into a latent representation $z=\mathcal{E}(x)$, and the decoder $D$ reconstructs the image from the latent, giving $x^{\prime}=D(z)=D(\mathcal{E}(x))$, where $z\in\mathbb{R}^{h\times w\times c}$. A domain-specific encoder $\tau_{\theta}$ projects $y$ (language prompt) to an intermediate representation $\tau_{\theta}(y)\in\mathbb{R}^{M\times d_{\tau}}$, which is then mapped to the intermediate layers of the U-Net (Ronneberger, Fischer, and Brox 2015) via a cross-attention layer implementing to condition the latent $z$. Based on image-conditioning pairs, the conditional LDM is learned via Equation 2.
@@
However when employing prompts like “a photo of a [occupation]”, the model has been shown to exhibit significant gender bias with stereotypes (Friedrich et al. 2023; Chuang et al. 2023), which manifests throughout the entire process.
While research efforts have been dedicated to addressing biases in the iterative image denoising process within U-Net, the convolutional network component of Stable Diffusion (Esposito et al. 2023), our investigation has revealed a distinct bias emerging after the text embedding stage when utilizing the text encoder of different versions of Stable Diffusion. This means that bias is amplified throughout the process.
We show a visualization of the dimensionality-reduced embeddings encoded by the text encoder of two versions of Stable Diffusion [^1] https://huggingface.co/runwayml/stable-diffusion-v1-5  [^2] https://huggingface.co/stabilityai/stable-diffusion-2-1-base using T-SNE, with prompts like “The photo of the face of a [occupation]”, where the [occupation] are from the top 8 male-biased and top 8 female-biased occupations from Fair Diffusion  [^3] https://github.com/ml-research/Fair-Diffusion/blob/main/results˙fairface˙generated˙1-5.txt, (details in Appendix B). There is a clear boundary between two gender-biases occupation encoded embeddings in both versions of Stable Diffusion. In other words, embeddings show gender biases regardless of whether they are encoded by CLIP or OpenCLIP (Radford et al. 2021; Schuhmann et al. 2022; Cherti et al. 2023), indicating that the text encoder already contains gender biases, which then conditions the U-Net. Thus, it is not enough to mitigate biases in the latent space of the U-Net, we need to address the text-encoded bias as well.
@@ 3.2 Measuring Text-Encoded Bias
Although DebiasVL claims great success in measuring bias in text embeddings and projecting it to achieve fairness, we have identified an issue. When attempting to switch the original prompt into another direction to achieve balance, the projection may take it into another direction that can introduce other biases. We identified failure cases in gender mitigation using DebiasVL’s method, which resulted in only male faces being generated, details in the Appendix F.
@@
We conduct zero-shot and unsupervised classification on embeddings, using prompts like “The photo of the face of a [occupation]” and a pre-trained Stable Diffusion text encoder. Our goal is to measure gender bias within these prompts and their embeddings. Altering the weights of the pre-trained text encoder is risky, as it could lead to the model forgetting information that is not related to gender, which has been learned from extensive data, so the Stable Diffusion text encoder is frozen. Another challenge is the lack of clear labels, as we only have prompts, embeddings, and statistics for each occupation generated by Fair Diffusion. These statistics are based on 250 images from the original Stable Diffusion and cannot be quantitatively analyzed or treated as a simple binary classification task due to the small sample size. Moreover, the embeddings are subject to change based on the hyperparameters of the pre-trained model, meaning the statistics can only serve as a benchmark for testing our approach rather than for model learning. Therefore, our analysis relies solely on prompts and model weights. As shown in Equation 1, $z$ is the prompt embedding, and $P$ is the projection process. Prompts $z_{0}$ applying Calibration Matrix can be written as $Pz_{0}=P_{0}\left(I+\frac{\lambda}{\left|S\right|}\sum_{(i,j)\in S}(z_{i}-z_{j})(z_{i}-z_{j})^{T}\right)^{-1}z_{0}$, derived from Equation 1. Thus, we can solve for $P$.
@@
$\Delta S(z_{0},z_{t},P_{0},P)$ $=\bigl{|}\text{Similarity}(P_{0}z_{0},P_{0}z_{t})$
$\quad-\text{Similarity}(Pz_{0},P_{0}z_{t})\bigr{|}$ (3)
@@
$\mathscr{G}(z_{0})$ $=\Delta S(z_{0},z_{male},P_{0},P)$
$\quad-\Delta S(z_{0},z_{female},P_{0},P)$ (4)
$z_{male}$ is the prompt with ‘male’ forcement, while $z_{female}$ is the prompt with ‘female’ forcement. For instance, $z_{0}$ is ‘a photo of the face of a nurse’, $z_{male}$ is ‘a photo of the face of a male nurse’, and $z_{female}$ is ‘a photo of the face of a female nurse’.
We assume the gender skew is male when $\mathscr{G}(z_{0})>0$ while the gender skew is female $\mathscr{G}(z_{0})<0$. If $\Delta S(z_{0},z_{t},P_{0},P)$ is larger, then the disparity between two prompts before and after applying the Calibration Matrix is also larger. This results in a greater divergence between the original embedding direction and the projected embeddings. Consequently, the strength of the redirected embeddings is increased, indicating a stronger gender preference before redirecting. In essence, if the redirection strength is substantial, it indicates a more forceful adjustment away from the original gender direction, which means a higher gender skew.
@@ 3.3 Defining the Bias Identification Gate
We utilize the Pearson Correlation Coefficient to measure similarity and check statistical approximation of genders in Fair Diffusion occupations, with further details provided in Appendix G. If the male count for a particular occupation surpasses half of the total, we assume this occupation is male-skewed; otherwise, it is female-skewed. To assess the effectiveness of our approach when designing the Bias Identification Gate, we compare the frequency-based labels with our calculations derived from Equation 3.2 to compute accuracy. We achieve an accuracy rate of 79%, indicating that our bias measurement approximation, achieved through task reframing, aligns with both intuition and statistical trends. The current accuracy is considered sufficient for mitigation since it is not a binary classification task. Additionally, we conducted extra experiments on skin tone after submission, achieving a 68% accuracy with the gating mechanism, and we can also be able to mitigate the bias.
The accuracy is lower when it comes to occupations such as insurance agents (details in Appendix I.1), as these are mostly likely at the boundary between male and female. In Table 1, we show mitigation of biases even when classifications of these challenging occupations are incorrect.
@@ 3.4 BiAs Experts: Bias Adapter Experts
The intuition is to guide the model to generate more female images when the original embedding exhibits the male skew and vice versa. To achieve this, we divide our training text-image pairs into two groups: male and female. We then fine-tune the model separately to generate male-biased and female-biased experts. When the gate is activated, it guides the experts according to the following rules: male skew for female-biased experts and female skew for male-biased experts. This process is called bias fine-tuning.
Figure 1: The Architecture of BiAs. We apply BiAs to the cross-attention of the U-Net. In the figure, the left BiAs represents the female expert, the right one, the male expert, and the middle one, the original cross-attention. The conditional information is processed by the Bias Identification Gate to determine which experts to choose - either male, female, or none of them. In the end, the chosen BiAs will process the input and be added together with the original cross-attention. The detailed BiAs architecture includes BiAs of the Q, K, V, and output matrices (for simplicity, we omit them in the figure). The parameters in cross-attention are frozen during the parameter-efficient fine-tuning process.
Adapters (Hu et al. 2022; Houlsby et al. 2019; Karimi Mahabadi, Henderson, and Ruder 2021) provide a method to freeze the model and introduce a new, trainable weight matrix, which significantly reduces both the time and memory required for training.
We incorporate this method into creating bias experts, making the bias fine-tuning parameter-efficient. We freeze the parameters in the U-Net and add adapters on the cross-attention layers of the U-Net, as illustrated in Figure 1. We initialize the adapters following the principles of LoRA (Hu et al. 2022): the weight parameters of the first matrix $W_{\text{down}}^{\prime}$ are determined by a Gaussian function, while the parameters of the second matrix $W_{\text{up}}^{\prime}$ are initialized as a zero matrix. Consequently, the added pathway is initially zero during training and does not impact the result, resembling the original output.
Since the adapters are randomly initialized, they do not inherit biases from Stable Diffusion fine-tuning, which makes them more effective in achieving fairness, as detailed in Section 4.2. By using adapters as the bias experts, we achieve better fairness scores and reduce the trainable ratio to only 5.6%.
For bias fine-tuning, we generate biased images with specific male or female characteristics so that the bias expert will learn stereotypes from the data. We employ a technique from Dreambooth (Ruiz et al. 2023), utilizing a special token to help BiAs to better understand the biased data. Dreambooth adds a special token, a unique identifier, to the prompt, and uses a few fine-tuning images of a subject as input to “personalize” the model. In our work, we use the special token to “personalize” the adapters of the biased information.
@@
For fine-tuning data, we generate our training image-text pairs using Stable Diffusion-XL with the following prompt: “A [gender] + [race] + [occupation].” The prompts generate 1530 images with different genders, races and, occupations to ensure variety, which then fine-tune the model using the special token. The results demonstrate that the special token yields better performance than ordinary fine-tuning, as detailed in Section 4.2.
@@
What distinguishes our work from methods such as Fair Diffusion is that we do not manually engage with gender attributes in the prompt or latent space; in other words, we do not introduce human intervention. The special token we add can be any token. Moreover, even without the special token, our approach still mitigates gender bias.